# Offline Model-Based Optimization: Comprehensive Review

### v1

@@ Abstract
Offline optimization is a fundamental challenge in science and engineering, where the goal is to optimize black-box functions using only offline datasets. This setting is particularly relevant when querying the objective function is prohibitively expensive or infeasible, with applications spanning protein engineering, material discovery, neural architecture search, and beyond. The main difficulty lies in accurately estimating the objective landscape beyond the available data, where extrapolations are fraught with significant epistemic uncertainty. This uncertainty can lead to objective hacking (reward hacking)—exploiting model inaccuracies in unseen regions—or other spurious optimizations that yield misleadingly high performance estimates outside the training distribution. Recent advances in model-based optimization (MBO) have harnessed the generalization capabilities of deep neural networks to develop offline-specific surrogate and generative models. Trained with carefully designed strategies, these models are more robust against out-of-distribution issues, facilitating the discovery of improved designs. Despite its growing impact in accelerating scientific discovery, the field lacks a comprehensive review. To bridge this gap, we present the first thorough review of offline MBO. We begin by formalizing the problem for both single-objective and multi-objective settings and by reviewing recent benchmarks and evaluation metrics. We then categorize existing approaches into two key areas: surrogate modeling, which emphasizes accurate function approximation in out-of-distribution regions, and generative modeling, which explores high-dimensional design spaces to identify high-performing designs. Finally, we examine the key challenges and propose promising directions for advancement in this rapidly evolving field including safe control of superintelligent systems. For a curated list of resources, please visit our repository.
@@ 1 Introduction
Offline optimization is a fundamental challenge in science and engineering, where the objective is to optimize a black-box function using only a fixed dataset (Trabucco et al., 2022). This setting has broad applications, including protein engineering (Sarkisyan et al., 2016b), material discovery (Hamidieh, 2018), and neural architecture search (Lu et al., 2023). For instance, in neural architecture search, the goal is to identify high-performing architectures solely from existing architecture-performance pairs, without training any new models which can be expensive. Unlike online optimization, which allows direct interaction with the objective function, offline optimization is particularly relevant when querying the function is costly, time-consuming, or infeasible (Angermüller et al., 2020; Barrera et al., 2016b; Sample et al., 2019b).
@@
Offline optimization is challenging because it requires accurately estimating the landscape of the black-box function beyond the available offline data (Trabucco et al., 2021). Extrapolating into these unseen regions suffers from significant epistemic uncertainty in two problematic scenarios. On the one hand, it may trigger reward hacking (Skalse et al., 2022) (which we refer to as objective hacking in our text), wherein the model exploits inaccuracies within the objective estimation in regions beyond training data. On the other hand, it can give rise to other forms of spurious optimization – especially in guided generative modeling – that yield misleadingly high performance estimates outside the training distribution (Brookes et al., 2019).
@@
Recent advances in offline model-based optimization (MBO) have harnessed the generalization power of deep neural networks to develop offline-specific surrogate and generative models. This progress has spurred two complementary lines of research. One line focuses on building surrogate models that extrapolate beyond the offline dataset, enabling robust function approximation and reliable gradient-based optimization to improve existing designs (Trabucco et al., 2021; Fu & Levine, 2021). The other line explores the use of generative models to navigate high-dimensional design spaces more effectively, facilitating the discovery of high-performing designs underrepresented in the offline data (Kumar & Levine, 2020; Kim et al., 2024b). Importantly, these two lines are not mutually exclusive – surrogate and generative models often complement each other to enhance overall performance (Fannjiang & Listgarten, 2020; Chen et al., 2024).
@@
Despite the rapid progress in offline MBO, both newcomers and seasoned researchers find it challenging to stay abreast of its evolving methodologies. Furthermore, the diversity of approaches and objectives has led to a fragmented landscape, making it difficult to discern overarching trends. To address these challenges, we present the first comprehensive review on offline MBO, synthesizing recent advances, categorizing key areas, and highlighting emerging directions. This review serves as both an accessible introduction for newcomers and a structured synthesis for experts looking to navigate the evolving frontiers of offline optimization.
@@
In this work, we formalize the problem settings for both offline single- and multi-objective optimization (Section 2). We also introduce a generative modeling perspective that frames offline optimization as conditional generation, compare single- and multi-objective settings, and discuss the connections between online and offline optimization. Additionally, we review recent benchmarks and propose a taxonomy that categorizes them into four application areas (Section 3): (1) synthetic function, (2) real-world system, (3) scientific design, and (4) machine learning (ML) model. Evaluation costs tend to increase – and our understanding of the underlying mechanisms tends to decrease – from categories (1) through (3). We address category (4) separately, given its growing prominence in the ML community. For each category, we detail the associated tasks, including the number of objectives and the oracle evaluators. We also provide an overview of the commonly used evaluation metrics, including usefulness, novelty, and diversity.
@@
Next, we categorize existing approaches into two key research lines—as we have discussed above: surrogate modeling (Section 4) and generative modeling (Section 5). Importantly, these approaches are not mutually exclusive; we explore in detail how they interact and complement one another. Finally, we conclude our paper (Section 6) by outlining promising future directions in this rapidly evolving field. In particular, we highlight several key areas for further exploration: (1) robust and realistic benchmarking, (2) uncertainty estimation of surrogate models, (3) causal graphical surrogate models, (4) advanced generative modeling, (5) and application to LLM alignment and AI Safety. An outline of our paper organization is depicted in Figure 1.
@@ 2 Problem Definition
In offline optimization, the goal is to discover a new design, denoted by $\bm{x}^{*}$, that maximizes the objective(s) $\bm{f}(\bm{x})$. This is achieved using an offline dataset $\mathcal{D}$, which consists of $N$ designs paired with their property labels. In particular, the dataset is given by
$\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}\,.$ (1)
where each design vector $\bm{x}_{i}$ belongs to a design space $\mathcal{X}\subseteq\mathbb{R}^{d}$, and each property label $\bm{y}_{i}\in\mathbb{R}^{m}$ contains the corresponding $m$ objective values for that design. The function $\bm{f}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ maps a design to its $m$-dimensional objective value vector.
@@
A critical challenge of this method is the accurate estimation of the objective landscape beyond the region covered by the training data. In these extrapolated regions, the surrogate model’s predictions are subject to significant epistemic uncertainty, potentially leading to objective (or reward) hacking. In other words, the model might exploit inaccuracies in regions lacking training data. We discuss strategies for building robust surrogates in Section 4.
@@
Find $\text{Find }\bm{x}^{*}\in\mathcal{X}\text{ such that there is no }\bm{x}\in\mathcal{X}\text{ with }\bm{f}(\bm{x})\succ\bm{f}(\bm{x}^{*})\,,$ (4)
@@
In addition to the surrogate modeling discussed above, generative modeling is another key ingredient in offline optimization. In fact, offline optimization methods can be viewed through the lens of conditional generation, where the objective is to model the distribution $p(\bm{x}\mid\bm{y}_{c})$ with $\bm{y}_{c}$ representing the desired conditions. By applying Bayes’ rule, this distribution can be decomposed as
@@
From this perspective, offline optimization methods generally fall into two categories: inverse and forward. Inverse methods directly train a conditional generative model for $p(\bm{x}\mid\bm{y}_{c})$. For example, MIN (Brookes et al., 2019) employs a GAN-based inverse mapping from $\bm{y}$ to $\bm{x}$ to generate designs that meet the desired specifications. Forward methods, on the other hand, leverage a surrogate model for $p(\bm{y}_{c}\mid\bm{x})$ to guide an unconditional generative model $p(\bm{x})$. For instance, ROMA (Yu et al., 2021) computes gradients in the latent space of a VAE to iteratively refine generated designs. Similarly, gradient-based approaches such as COMs (Trabucco et al., 2021) and ICT (Yuan et al., 2023) fall into this category, although many of these methods do not explicitly model the generative component $p(\bm{x})$.
Since the literature on offline optimization often integrates surrogate and generative modeling, we separate their discussion in this review to provide clearer insights into each component: surrogate modeling is discussed in Section 4, while generative modeling is covered in Section 5.
@@
Another fundamental difference is the sampling strategy. In SOO, the focus is on a single property, which simplifies conditional generation to either computing the gradient of that property (Chen et al., 2023a) or building an inverse mapping from the property to the design (Kumar & Levine, 2020). These approaches, however, are not directly applicable to MOO, where improvements in one objective might lead to the deterioration of another. Consequently, MOO relies on Pareto-aware sampling strategies to navigate the trade-offs among conflicting objectives. For example, Paretoflow (Yuan et al., 2024a) assigns uniform weight vectors to different objectives to guide flow matching towards the Pareto front.
@@
Relation with Online Optimization (Active Learning).
Online optimization (often referred to as active learning in this context) involves iteratively querying an expensive ground-truth objective function for new design points and then using these new data points to update a surrogate model (Jain et al., 2022; Gruver et al., 2023; Frey et al., 2025). In high-cost domains such as drug discovery—where each experiment (e.g., a clinical trial) can be prohibitively expensive (Angermüller et al., 2020)—these queries must be made sparingly under a limited budget. The crux of online optimization is to strategically select new points that are both likely to yield high rewards and carry substantial uncertainty, thereby maximizing information gain and accelerating the search.
Bayesian models, such as Gaussian Processes (GPs) (Williams & Rasmussen, 1995; Rasmussen & Williams, 2006), are commonly used for online optimization because they provide principled uncertainty quantification. However, GPs have cubic complexity in the number of data points, which becomes infeasible for large datasets. As a result, deep neural networks equipped with approximate uncertainty estimation techniques (e.g., Monte Carlo dropout or ensembles) are often employed, though their uncertainty estimates can be less reliable in practice (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017).
In offline optimization, there is no ability to query the objective after an initial dataset has been collected—effectively a zero-round version of active learning. Since no further data can be acquired, offline methods often avoid high-uncertainty regions where the surrogate model might be highly inaccurate or risky. Consequently, offline optimization algorithms often adopt more conservative strategies, prioritizing regions of the design space that are both high reward and well-understood according to the existing dataset.
@@
Futhermore online and offline approaches can synergize naturally. In the online phase, the goal is to collect a high-quality dataset by judiciously querying promising and uncertain regions. Once the budget for queries is exhausted and the dataset becomes static, offline optimization can then leverage the collected data for a final decision. In this way, online optimization focuses on acquiring the most informative dataset possible, while offline optimization focuses on extracting the best solution from the available data.
@@ 3.1 Task
It is important to note that these categories are not mutually exclusive; some benchmarks may belong to more than one category. For example, Ehrlich (Stanton et al., 2024) can be viewed as a synthetic function task for biological sequence optimization, yet we mainly discuss it under scientific design to better reflect its application context. Each subsubsection first discusses the intrinsic advantages and limitations of its task category and then introduces the commonly used tasks within that category. For each task, we report its size (i.e., the offline dataset), space (design space), dimension (design dimension), # Obj (number of objectives), and oracle (evaluation oracle type), as summarized in Tables 1, 2, 3, and 4. This organization offers a balanced view of both the theoretical appeal and practical applicability of each task category.

### v2

@@ Abstract
Offline black-box optimization is a fundamental challenge in science and engineering, where the goal is to optimize black-box functions using only offline datasets. This setting is particularly relevant when querying the objective function is prohibitively expensive or infeasible, with applications spanning protein engineering, material discovery, neural architecture search, and beyond. The main difficulty lies in accurately estimating the objective landscape beyond the available data, where extrapolations are fraught with significant epistemic uncertainty. This uncertainty can lead to objective hacking (reward hacking)—exploiting model inaccuracies in unseen regions—or other spurious optimizations that yield misleadingly high performance estimates outside the offline distribution. Recent advances in model-based optimization (MBO) have harnessed the generalization capabilities of deep neural networks to develop offline-specific surrogate and generative models. Trained with carefully designed strategies, these models are more robust against out-of-distribution issues, facilitating the discovery of improved designs. Despite its growing impact in accelerating scientific discovery, the field lacks a comprehensive review. To bridge this gap, we present the first thorough review of offline MBO. We begin by formalizing the problem for both single-objective and multi-objective settings and by reviewing recent benchmarks and evaluation metrics. We then categorize existing approaches into two key areas: surrogate modeling, which emphasizes accurate function approximation in out-of-distribution regions, and generative modeling, which explores high-dimensional design spaces to identify high-performing designs. Finally, we examine the key challenges and propose promising directions for advancement in this rapidly evolving field including safe control of superintelligent systems. For a curated list of resources, please visit our repository.
@@ 1 Introduction
Offline black-box optimization is a fundamental challenge in science and engineering, where the objective is to optimize a black-box function using only a fixed dataset (Trabucco et al., 2022). This setting has broad applications, including protein engineering (Sarkisyan et al., 2016; Yang et al., 2019), material discovery (Hamidieh, 2018), and neural architecture search (Lu et al., 2023). For instance, in neural architecture search, the goal is to identify high-performing architectures solely from existing architecture-performance pairs, without training any new models which can be expensive. Unlike online black-box optimization, which allows direct interaction with the objective function, offline black-box optimization is particularly relevant when querying the function is costly, time-consuming, or infeasible (Angermüller et al., 2020; Barrera et al., 2016; Sample et al., 2019).
@@
Offline black-box optimization is challenging because it requires accurately estimating the landscape of the black-box function beyond the available offline data (Trabucco et al., 2021). Extrapolating into these unseen regions suffers from significant epistemic uncertainty, arising from limited data coverage and imperfect model knowledge rather than inherent noise (Der Kiureghian & Ditlevsen, 2009), which manifests in two problematic scenarios. On the one hand, it may trigger reward hacking (Skalse et al., 2022) (which we refer to as objective hacking in our text), wherein the model exploits inaccuracies within the objective estimation in regions beyond offline data. On the other hand, it can give rise to other forms of spurious optimization – especially in guided generative modeling where the model overfits spurious correlations within offline data – that yield misleadingly high-performance sample generation outside offline distribution (Brookes et al., 2019).
@@
Offline model-based optimization (MBO) is a paradigm that leverages the generalization capabilities of deep neural networks to develop offline-specific surrogate and generative models for solving offline black-box optimization problems. This progress has spurred two complementary lines of research. One line focuses on building surrogate models that extrapolate beyond the offline dataset, enabling robust function approximation and reliable gradient-based optimization to improve existing designs (Trabucco et al., 2021; Fu & Levine, 2021). The other line explores the use of generative models to navigate high-dimensional design spaces more effectively, facilitating the discovery of high-performing designs underrepresented in the offline data (Kumar & Levine, 2020; Kim et al., 2024b). Importantly, these two lines are not mutually exclusive – surrogate and generative models often complement each other to enhance overall performance (Fannjiang & Listgarten, 2020; Chen et al., 2024). The design-centric offline MBO differs from offline RL (Levine et al., 2020; Prudencio et al., 2023), which seeks to learn an optimal policy from trajectory data and emphasizes sequential decision-making and credit assignment in Markov decision processes, whereas offline MBO focuses on single-step design generation from static datasets.
@@
Despite the rapid progress in offline MBO, both newcomers and seasoned researchers find it challenging to stay abreast of its evolving methodologies. Furthermore, the diversity of approaches and objectives has led to a fragmented landscape, making it difficult to discern overarching trends. To address these challenges, we present the first comprehensive review on offline MBO, synthesizing recent advances, categorizing key areas, and highlighting emerging directions. This review serves as both an accessible introduction for newcomers and a structured synthesis for experts looking to navigate the evolving frontiers of offline MBO.
@@
In this work, we formalize the problem settings for both offline single- and multi-objective optimization (Section 2). We also introduce a generative modeling perspective that frames offline black-box optimization as conditional generation, compare single- and multi-objective settings, and discuss the connections between online and offline black-box optimization. Additionally, we review recent benchmarks and propose a taxonomy that categorizes them into four application areas (Section 3): (1) synthetic function, (2) real-world system, (3) scientific design, and (4) machine learning (ML) model. Evaluation costs tend to increase – and our understanding of the underlying mechanisms tends to decrease – from categories (1) through (3). We address category (4) separately, given its growing prominence in the ML community. For each category, we detail the associated tasks, including the number of objectives and the oracle evaluators. We also provide an overview of the commonly used evaluation metrics, including usefulness, novelty, and diversity.
@@
Next, we categorize existing approaches into two key research lines—as we have discussed above: surrogate modeling (Section 4) and generative modeling (Section 5). Importantly, these approaches are not mutually exclusive, and surrogate models and generative models are often used together in offline MBO. We separate their discussion in this review to provide clearer insights into each component, while also exploring in detail how they interact and complement one another. Finally, we conclude our paper (Section 6) by outlining promising future directions in this rapidly evolving field. In particular, we highlight several key areas for further exploration: (1) robust and realistic benchmarking, (2) uncertainty estimation of surrogate models, (3) causal graphical surrogate models, (4) advanced generative modeling, (5) and application to LLM alignment and AI Safety. An outline of our paper organization is depicted in Figure 1.
@@ 2 Problem Definition
In offline black-box optimization, the goal is to discover a new design, denoted by $\bm{x}^{*}$, that maximizes the objective(s) $\bm{f}(\bm{x})$. This is achieved using an offline dataset $\mathcal{D}$, which consists of $N$ designs paired with their property labels. In particular, the dataset is given by
$\mathcal{D}=\{(\bm{x}_{i},\bm{y}_{i})\}_{i=1}^{N}$ (1)
where each design vector $\bm{x}_{i}$ belongs to a design space $\mathcal{X}\subseteq\mathbb{R}^{d}$, and each property label $\bm{y}_{i}\in\mathbb{R}^{m}$ contains the corresponding $m$ objective values for that design. The function $\bm{f}:\mathcal{X}\rightarrow\mathbb{R}^{m}$ maps a design to its $m$-dimensional objective value vector. Offline MBO applies to static optimization problems where the objective $\bm{f}(\bm{x})$ is a fixed black-box function without temporal structure, distinguishing it from sequential decision-making problems such as system identification followed by optimal control (Sutton & Barto, 1998).
@@
A critical challenge of this method is the accurate estimation of the objective landscape beyond the region covered by the offline data. In these extrapolated regions, the surrogate’s posterior predictive variance captures its epistemic uncertainty—the component due to limited data and model knowledge rather than inherent noise. Formally, if we place a parameter posterior $p(\bm{\theta}\mid\mathcal{D})$ over model parameters, then the predictive distribution under this posterior is
$p(y\mid\bm{x},\mathcal{D})=\int p(y\mid\bm{x},\bm{\theta})\,p(\bm{\theta}\mid\mathcal{D})\,d\bm{\theta},$
whose total variance decomposes into
$\mathrm{Var}[y\mid\bm{x},\mathcal{D}]=\underbrace{\mathbb{E}_{p(\bm{\theta}\mid\mathcal{D})}\bigl[\mathrm{Var}(y\mid\bm{x},\bm{\theta})\bigr]}_{\text{aleatoric}}\;+\;\underbrace{\mathrm{Var}_{p(\bm{\theta}\mid\mathcal{D})}\bigl[\mathbb{E}(y\mid\bm{x},\bm{\theta})\bigr]}_{\text{epistemic}},$
where the first term quantifies aleatoric uncertainty arising from inherent noise in the data, and the second term quantifies epistemic uncertainty and diminishes as more data are observed (Der Kiureghian & Ditlevsen, 2009). In practice, epistemic uncertainty can be estimated via:
  • Gaussian Processes: posterior variance gives closed-form decomposition (Rasmussen & Williams, 2006).
  • Deep Ensembles: variance across multiple network predictions approximates parameter uncertainty (Lakshminarayanan et al., 2017).
  • Monte Carlo Dropout: repeated stochastic forward passes approximate the weight posterior (Gal & Ghahramani, 2016).
  • Bayesian Neural Networks: variational or sampling-based inference yields predictive distributions with parameter uncertainty.
High epistemic uncertainty in out-of-distribution regions can lead to objective hacking, where the optimizer exploits surrogate inaccuracies to identify spurious high-performing designs. We discuss strategies for building robust surrogates in Section 4.
@@
[…]
@@
In addition to the surrogate modeling discussed above, generative modeling is another key ingredient in offline MBO. In fact, offline MBO methods can be viewed through the lens of conditional generation, where the objective is to model the distribution $p(\bm{x}\mid\bm{y}_{c})$ with $\bm{y}_{c}$ representing the desired conditions. By applying Bayes’ rule, this distribution can be decomposed as
@@
From this perspective, offline MBO methods generally fall into two categories: inverse and forward. Inverse methods directly train a conditional generative model for $p(\bm{x}\mid\bm{y}_{c})$. For example, MIN (Brookes et al., 2019) employs the generative adversarial network (GAN)-based inverse mapping from $\bm{y}$ to $\bm{x}$ to generate designs that meet the desired specifications. Forward methods, on the other hand, leverage a surrogate model for $p(\bm{y}_{c}\mid\bm{x})$ to guide an unconditional generative model $p(\bm{x})$. For instance, ROMA (Yu et al., 2021a) computes gradients in the latent space of a variational autoencoder (VAE) to iteratively refine generated designs. Similarly, gradient-based approaches such as conservative objective models (COMs) (Trabucco et al., 2021) and ICT (Yuan et al., 2023) fall into this category, although many of these methods do not explicitly model the generative component $p(\bm{x})$.
Since the literature on offline MBO often integrates surrogate and generative modeling, we separate their discussion in this review to provide clearer insights into each component: surrogate modeling is discussed in Section 4, while generative modeling is covered in Section 5. We discuss generative modeling here only at a high level to provide a novel conditional-generation view of offline MBO, and revisit it in Section 5 for a detailed exposition of methodological developments.
@@
Another fundamental difference is the sampling strategy. In SOO, the focus is on a single property–such as model accuracy in neural architecture design–which simplifies conditional generation to either computing the gradient of that property (Chen et al., 2023a) or building an inverse mapping from the property to the design (Kumar & Levine, 2020). These approaches, however, are not directly applicable to MOO, where improvements in one objective (e.g., accuracy) might lead to the deterioration of another (e.g., efficiency). Consequently, MOO relies on Pareto-aware sampling strategies to navigate the trade-offs among conflicting objectives. For example, Paretoflow (Yuan et al., 2025) assigns uniform weight vectors to different objectives to guide flow matching towards the Pareto front.
@@
Relation with Online Black-Box Optimization.
online black-box optimization involves iteratively querying an expensive ground-truth objective function for new design points and then using these new data points to update a surrogate model (Jain et al., 2022; Gruver et al., 2023; Frey et al., 2025). In high-cost domains such as drug discovery—where each experiment (e.g., a clinical trial) can be prohibitively expensive (Angermüller et al., 2020)—these queries must be made sparingly under a limited budget. A central challenge in online black-box optimization is balancing the trade-off between exploitation—selecting designs with high predicted objective values—and exploration—querying uncertain regions to improve the model’s understanding of the objective landscape. This exploration-exploitation dilemma has been widely studied in the context of Bayesian optimization (Frazier, 2018), where acquisition functions such as the Upper Confidence Bound and Expected Improvement are commonly used to guide query selection. This principle has also been explored in the operations research, notably through Efficient Global Optimization (EGO) (Jones et al., 1998).
Bayesian models, such as Gaussian Processes (GPs) (Williams & Rasmussen, 1995; Rasmussen & Williams, 2006), are commonly used for online black-box optimization because they provide principled uncertainty quantification. However, GPs have cubic complexity in the number of data points, which becomes infeasible for large datasets. As a result, deep neural networks equipped with approximate uncertainty estimation techniques (e.g., Monte Carlo dropout or ensembles) are often employed, though their uncertainty estimates can be less reliable in practice (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017).
In offline black-box optimization, there is no ability to query the objective after an initial dataset has been collected—effectively a zero-round version of online black-box optimization. Since no further data can be acquired, offline methods often avoid high-uncertainty regions where the surrogate model might be highly inaccurate or risky. Consequently, offline black-box optimization algorithms often adopt more conservative strategies, prioritizing regions of the design space that are both high reward and well-understood according to the existing dataset. One example from online black-box optimization is the use of the Upper Confidence Bound (Frazier, 2018), an acquisition function that promotes exploration by combining predicted reward with uncertainty. By adopting a Lower Confidence Bound (Frazier, 2018), another acquisition function that promotes exploitation and favors high-reward but low-uncertainty regions, the same principle can be adapted for offline MBO. This contrast illustrates the core difference between online and offline black-box optimization: exploration versus conservatism.
@@
Futhermore online and offline methods can synergize naturally. In the online phase, the goal is to collect a high-quality dataset by judiciously querying promising and uncertain regions. Once the budget for queries is exhausted and the dataset becomes static, offline black-box optimization can then leverage the collected data for a final decision. In this way, online black-box optimization focuses on acquiring the most informative dataset, while offline black-box optimization focuses on extracting the best solution from the available data.
@@ 3.1 Task
It is important to note that these categories are not mutually exclusive; some benchmarks may belong to more than one category. For example, Ehrlich (Stanton et al., 2024) can be viewed as a synthetic function task for biological sequence optimization, yet we mainly discuss it under scientific design to better reflect its application context. Each subsubsection first discusses the intrinsic advantages and limitations of its task category and then introduces the commonly used tasks within that category. For each task, we report its size (i.e., the offline dataset), space (design space), dimension (design dimension), # Obj (number of objectives), and oracle (evaluation oracle type). This organization offers a balanced view of both the theoretical appeal and practical applicability of each task category.